At 7:29 a.m. the temperature was -4 degrees F; at 7:32 a.m. the temperature was 45 degrees F. What was the average rate of change in temperature (per minute)?
A) 16.3 B) 49 C) -16.3 D) 24.5

Answer by checkley77(12569)

7:32-7:29=:03 or 3 minutes.
45-(-4)=45+4=49 degree change of temperature.
49/3=16.3 deg. per minute.